# UGM2N: An Unsupervised and Generalizable Mesh Movement Network via M-Uniform Loss

### v1

@@ 2 Related Work
For mesh movement method, Omella and Pardo [30] proposed a neural network-enhanced boundary node optimization method, which is specifically designed for tensor product meshes. Song et al. [9] introduced M2N, a framework combining neural splines with GAT, enabling end-to-end mesh movement with 3–4 order-of-magnitude speedups. Hu et al. [31] introduced a neural mesh adapter trained via the MA equation physical loss to dynamically adjust mesh nodes, and develops a moving mesh neural PDE solver that improves modeling accuracy for dynamic systems. Rowbottom et al. [32] proposed a graph neural diffusion method that directly minimize finite element error to achieve efficient mesh adaptation. For specialized applications, methods such as Flow2Mesh and Para2Mesh have demonstrated the efficacy of learning-based adaptation in aerodynamic simulations [33, 34]. Recent work extended these advances with UM2N [10], a universal graph-transformer architecture attempts to achieve zero-shot adaptation across diverse PDEs and geometries. Most of the aforementioned works rely on supervised learning, where models are trained to align their outputs with pre-adapted meshes, resulting in a lack of physical information. Additionally, they often require retraining for different PDEs or mesh geometries, limiting their generalizability. This paper adopts an unsupervised learning approach to achieve equation-agnostic generalization across arbitrary mesh geometries.
@@ 3.1 Problem statement and preliminaries
Given an initial mesh $\mathcal{M}$ (e.g., a uniform mesh) and associated flow field variables (such as velocity $\mathbf{u}$ or pressure $p$), the mesh movement method optimizes the node positions to generate an adapted mesh satisfying predefined resolution criteria. The mesh movement process can be analyzed from different perspectives, such as coordinate mapping between uniform and adapted meshes, uniform mesh construction in metric space, and so on [6]. The MA-based method adopts the former approach, solving the MA equation with boundary conditions to obtain the coordinate mapping. In contrast, this study employs the latter perspective, enforcing uniform distribution in metric space without explicit coordinate transformations between computational and physical domains.
@@
where $\sigma=\int_{\Omega}\rho(\mathbf{x})\,d\mathbf{x}$, $m(\mathbf{x})=\sqrt{\det(M(\mathbf{x}))}$ is the mesh density function, $K$ is the element of $\mathcal{M}$, and $N_{e}$ is the number of elements in the mesh $\mathcal{M}$. This condition constrains the size of mesh elements—when $m(\mathbf{x})$ is large, the element volume should be small, and vice versa. Additionally, the M-Uniform mesh condition also requires that the mesh elements should be equilateral in the metric space. In this work, we primarily focus on modifying the mesh density while disregarding the equilateral alignment of mesh elements.
@@ 3.2 Network overview
Iterative mesh adaptation with dynamic termination. Inspired by iterative mesh smoothing, multi-epoch mesh adaptation is employed to progressively refine node distribution during inference [^1] Multi-epoch mesh adaptation is disabled during training to simplify the training process., the Hessian norm values are updated between epochs via Delaunay triangulation-based interpolation from original to adapted nodes, providing initialization for subsequent adaptations. While this process could theoretically continue indefinitely, convergence is not guaranteed and mesh validity may degrade. A fixed epoch count would limit optimization capability; instead, we propose a metric-based adaptive strategy that dynamically determines termination based on optimization progress. This approach will be detailed following the presentation of our unsupervised loss function.
@@ 3.3 M-Uniform loss
where $\mathbf{\theta}$ is the model parameters, and $\lambda=100$ is a scaling constant. This approach shares conceptual similarities with PINNs, where local constraints—residual conditions in PINNs and the M-Uniform condition here—guide the learning process. By enforcing mesh equidistribution condition at the node level, the model can adapt mesh node positions without supervised data, ensuring generalization to arbitrary adaptive scenarios. Crucially, unlike existing adaptive methods (e.g., M2N or UM2N), training does not require the full mesh as input. Instead, it can be trained on individual mesh nodes. For example, during mini-batch training, we can sample a subset of mesh nodes from the mesh and achieve efficient training through graph batching. Moreover, this approach further reduces the required amount of data, as the number of data samples is proportional to the number of mesh nodes rather than the number of meshes.
@@
where $N_{e}$ is the number of mesh elements in $\mathcal{M^{\prime}}$, and $\overline{\mathcal{L}_{K_{l}}}=\frac{1}{N_{e}}\sum_{l=1}^{N_{e}}\mathcal{L}_{K_{l}}$. The model stops the iterative mesh adaptation when $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ no longer decreases. During inference, we set a fixed upper limit for the number of iterations—specifically, we set the maximum number of mesh adaptation epochs as 10. The full adaptation algorithm is detailed in App. A.1, and the theoretical analysis of the effectiveness of the loss function to optimize mesh distribution is provided in App. A.2.
@@ 4.1 Experiment setups
Model training. Following UM2N’s protocol, we trained UGM2N on a mesh with only four flow fields and evaluated its zero-shot generalization performance on unseen flow fields or meshes. As depicted in Fig. 10 (App. B.1), random perturbations were applied to mesh node positions to enhance data diversity, resulting in a training set comprising 10,440 mesh nodes. The model was optimized using Nadam [36] with an initial learning rate of 1e-4, with all experiments conducted on an NVIDIA RTX TITAN GPU. See App. C for more training details.
Baselines and metrics. We performed a comparative analysis against the MA method [37], M2N, and UM2N, using UM2N’s pre-trained weights obtained from its GitHub repository [38]. Performance was evaluated using two key metrics: (1) error reduction (ER), which measures the relative improvement in PDE solution accuracy compared to the initial coarse mesh (with the high-resolution solution serving as the reference), and (2) tangling ratio (TR), defined as the fraction of invalid elements in the adapted mesh. Additional case-specific metrics will be presented in the corresponding experimental sections. Detailed mathematical definitions of all metrics are provided in App. F.
@@ 4.2 Performances across different flow field solutions
To assess the model’s generalization ability across diverse flow fields, we conducted experiments using Burgers’ equation with varying initial conditions, as well as Poisson and Helmholtz equations with different analytical solutions (refer to App. B.2 and B.3 for more detailed PDE configurations). All simulations were performed on a uniform triangular mesh spanning the domain $0,1]\times[0,1]$, comprising 1,478 elements. For validation, a high-fidelity reference solution was computed on a significantly refined mesh with 23,250 elements, ensuring precise accuracy for benchmarking purposes.
The test results are summarized in Table 1, it can be observed that the our model demonstrates significant advantages in the vast majority of cases. In the seven test cases for the Poisson equation, our model achieved optimal performance (highest ER or lowest TR) in five cases, particularly excelling with complex functions. For example, for $\sum_{i,j}\exp[-\frac{(x-x_{\mu,i})^{2}}{0.25^{2}}-\frac{(y-y_{\mu,j})^{2}}{0.25^{2}}]$, ours achieved an ER of 9% , far surpassing the comparison methods (MA: -8.64%, M2N: -30.20%).
@@
Additionally, ours achieved complete dominance in the Helmholtz equation, securing the four highest ER out of five test cases. Although slightly inferior to M2N in the Burgers equation, ours still significantly outperformed UM2N. These results validate the robustness and generalization capability of our method in mesh adaptation across different PDEs, achieving state-of-the-art performance compared to existing approaches.
The mesh adaptation results for the Helmholtz equation are shown in Fig. 2 (for results on the Poisson variance and Burgers equation, refer to App. D). Our method demonstrates superior visual alignment with the target flow field while simultaneously generating meshes with excellent quality compared to alternative approaches. It is noteworthy that although M2N and UM2N can also produce approximately adaptive results, they show weaker adherence to the mesh equidistribution condition compared to the our method (see App. E).
@@ 4.3 Performance across varying mesh geometries
Figure 3: The ER for different mesh resolutions on the Helmholtz equation. For clarity in presentation, we clipped the minimum ER at -1%, even though for some methods (e.g., M2N), their adapted meshes significantly increased the solution error.
@@
Regarding computational efficiency, as illustrated in Fig. 4, we present the time required for mesh movement under varying element sizes. For each model configuration, we conducted ten repeated tests on different solutions of the Helmholtz equation and reported the average mesh movement time per trial. Compared with the MA method, the neural-based mesh movement method demonstrates significant advantages. Although the iterative mesh adaptation in UGM2N introduces computational overhead proportional to the optimization epoch $E$, its computational efficiency does not exhibit substantial degradation when compared to M2N or UM2N. The reason lies in its relatively low per-epoch optimization time (indicated by the dashed line in the figure).
Varying mesh shapes. Another potential variable in the simulation is the shape of mesh. To evaluate the model’s generalization capability under different mesh geometric configurations, we quantitatively conducted three distinct simulation cases: subsonic flow around a NACA0012 airfoil, cylinder flow, and the wave equation on a circular domain, with the experimental setups provided in App. B.2. The adaptive results of the airfoil mesh are shown in Fig. 6, which shows that our model effectively captures the shock wave location without introducing any invalid elements. Additionally, at the position $y=0.25$, our model obtains the minimum mean absolute error (MAE  [^3] The average absolute error between the solution on the high-resolution mesh and the solution on the adaptive (coarse) mesh at each position (time step).) in pressure coefficient result. For the cylinder flow case, neither MA nor M2N could produce valid meshes. As shown in Fig. 6, compared to UM2N, our model further reduces the prediction error of the drag coefficient and slightly decreasing the cumulative error during the solving process. The results of the wave equation are shown in Fig. 8. At any given time step, our proposed method consistently generates smooth, high-quality adaptive meshes. In contrast, other methods may produce distorted mesh elements and fail to reduce errors, although they can also generate adaptive meshes.
Figure 5: Results on the cylinder flow. We present the adaptive meshes at four time slices, with the results over the entire simulation period provided in the supplementary vedio materials.
@@
Figure 8: Mesh adaptation on moving cylinder and supersonic flow cases. In supersonic flow, UGM2N’s adaptive mesh dynamically adjusts nodes to track shock waves; in the moving cylinder case, it precisely captures both the cylinder boundary and wake flow. The results are also presented in the supplementary vedio.
In addition to the above quantitative experiments, two additional qualitative experiments—supersonic flow over a wedge and a moving cylinder inside a channel—are provided in Fig. 8; the results also demonstrate that our model can also generate effective adaptive meshes for more complex cases.
@@ 4.4 Ablation Study
Loss function To validate the effectiveness of the M-Uniform loss compared to the coordinate loss of M2N and the volume loss of UM2N, we trained models with the same architecture and the same training data but different loss functions. The supervised data was generated using the MA method. Table 9 presents the average ER on PDEs with different solutions of models trained with different loss functions on three types of equations (the equation configurations are the same as in Table 1). With only a small amount of training data, supervised learning methods using the entire mesh as input struggle to produce effective models. In contrast, our unsupervised training approach requires no adaptive meshes as supervised data, and the node patch-based training method enables effective model training even with limited data.
@@ 5 Conclusion
We introduce UGM2N, an unsupervised and generalizable mesh movement network. This network removes the requirement for pre-adapted meshes while demonstrating strong generalization capabilities across various PDEs, geometric configurations, and mesh resolutions. By leveraging localized node-patch representations and a novel M-Uniform loss, our approach enforces mesh equidistribution properties comparable to Monge-Ampère-based methods—but in a more efficient, unsupervised manner. Extensive experiments demonstrate consistent performance improvements over both supervised learning baselines and traditional mesh adaptation techniques, achieving significant error reductions without mesh tangling across diverse PDEs and mesh geometries. See App. G for the discussions on limitations and broader impacts.
Appendix A Method
A.1 The proposed mesh movement method

### v2

@@ 2 Related Work
For mesh movement method, Omella and Pardo [30] proposed a neural network-enhanced boundary node optimization method, which is specifically designed for tensor product meshes. Song et al. [9] introduced M2N, a framework combining neural splines with Graph Attention Network (GAT) [31], enabling end-to-end mesh movement with 3–4 order-of-magnitude speedups. Hu et al. [32] introduced a neural mesh adapter trained via the MA equation physical loss to dynamically adjust mesh nodes, and develops a moving mesh neural PDE solver that improves modeling accuracy for dynamic systems. Rowbottom et al. [33] proposed a graph neural diffusion method that directly minimizes finite element error to achieve efficient mesh adaptation, and proved that its model architecture can effectively avoid mesh entanglement. For specialized applications, methods such as Flow2Mesh and Para2Mesh have demonstrated the efficacy of learning-based adaptation in aerodynamic simulations [34, 35]. Recent work extended these advances with UM2N [10], a universal graph-transformer architecture attempts to achieve zero-shot adaptation across diverse PDEs and geometries. Most of the aforementioned works rely on supervised learning, where models are trained to align their outputs with pre-adapted meshes, resulting in a lack of physical information. Additionally, they often require retraining for different PDEs or mesh geometries, limiting their generalizability. This paper adopts an unsupervised learning approach to achieve equation-agnostic generalization across arbitrary mesh geometries.
@@ 3.1 Problem statement and preliminaries
Given an initial mesh $\mathcal{M}$ (e.g., a uniform mesh) and associated flow field variables (such as velocity $\mathbf{u}$ or pressure $p$), the mesh movement method optimizes the node positions to generate an adapted mesh satisfying predefined resolution criteria. The mesh movement process can be analyzed from different perspectives, such as coordinate mapping between uniform and adapted meshes, uniform mesh construction in metric space, and so on [6]. The MA-based method adopts the former approach, solving the MA equation with boundary conditions to obtain the coordinate mapping (for a detailed introduction, refer to App. A). In contrast, this study employs the latter perspective, enforcing uniform distribution in metric space without explicit coordinate transformations between computational and physical domains.
@@
where $\sigma=\int_{\Omega}m(\mathbf{x})\,d\mathbf{x}$, $m(\mathbf{x})=\sqrt{\det(M(\mathbf{x}))}$ is the mesh density function, $K$ is the element of $\mathcal{M}$, and $N_{e}$ is the number of elements in the mesh $\mathcal{M}$. This condition constrains the size of mesh elements—when $m(\mathbf{x})$ is large, the element volume should be small, and vice versa. Additionally, the M-Uniform mesh condition also requires that the mesh elements should be equilateral in the metric space. In this work, we primarily focus on modifying the mesh density while disregarding the equilateral alignment of mesh elements.
@@ 3.2 Network overview
Iterative mesh adaptation with dynamic termination. Inspired by iterative mesh smoothing, multi-epoch mesh adaptation is employed to progressively refine node distribution during inference [^1] Multi-epoch mesh adaptation is disabled during training to simplify the training process., the Hessian norm values are updated between epochs via Delaunay triangulation-based interpolation from original to adapted nodes, providing initialization for subsequent adaptations. While this process could theoretically continue indefinitely, convergence is not guaranteed and mesh validity may degrade (see App. B.2 for discussions). A fixed epoch count would limit optimization capability; instead, we propose a metric-based adaptive strategy that dynamically determines termination based on optimization progress. This approach will be detailed following the presentation of our unsupervised loss function.
@@ 3.3 M-Uniform loss
where $\mathbf{\theta}$ is the model parameters, and $\lambda=100$ is a scaling constant. This approach shares conceptual similarities with PINNs, where local constraints—residual conditions in PINNs and the M-Uniform condition here—guide the learning process. By enforcing mesh equidistribution condition at the node level, the model can adapt mesh node positions without supervised data, ensuring generalization to arbitrary adaptive scenarios. Crucially, unlike existing adaptive methods (e.g., M2N or UM2N), training does not require the full mesh as input. Instead, it can be trained on individual mesh nodes. For example, during mini-batch training, we can sample a subset of mesh nodes from the mesh and achieve efficient training through graph batching [^2] Notably, during training, we do not use the model outputs to update the mesh nodes—in other words, there is no dynamic mesh update.. Moreover, this approach further reduces the required amount of data, as the number of data samples is proportional to the number of mesh nodes rather than the number of meshes.
@@
where $N_{e}$ is the number of mesh elements in $\mathcal{M^{\prime}}$, and $\overline{\mathcal{L}_{K_{l}}}=\frac{1}{N_{e}}\sum_{l=1}^{N_{e}}\mathcal{L}_{K_{l}}$. The model stops the iterative mesh adaptation when $\mathcal{L}_{\mathrm{var}}\left(\mathcal{M^{\prime}}\right)$ no longer decreases. During inference, we set a fixed upper limit for the number of iterations—specifically, we set the maximum number of mesh adaptation epochs as 10. The full adaptation algorithm is detailed in App. B.1, and the theoretical analysis of the effectiveness of the loss function to optimize mesh distribution is provided in App. B.2.
@@ 4.1 Experiment setups
Model training. Following UM2N’s protocol, we trained UGM2N on a mesh with only four flow fields and evaluated its zero-shot generalization performance on unseen flow fields or meshes. As depicted in Fig. 12 (App. C.1), random perturbations were applied to mesh node positions to enhance data diversity, resulting in a training set comprising 10,440 mesh nodes. The model was optimized using Nadam [37] with an initial learning rate of 1e-4, with all experiments conducted on an NVIDIA RTX TITAN GPU. See App. D for more training details.
Baselines and metrics. We performed a comparative analysis against the MA method [38], M2N, and UM2N, using UM2N’s pre-trained weights obtained from its GitHub repository [39]. Performance was evaluated using two key metrics: (1) error reduction (ER), which measures the relative improvement in PDE solution accuracy compared to the initial coarse mesh (with the high-resolution solution serving as the reference), and (2) tangling ratio (TR), defined as the fraction of invalid elements in the adapted mesh. Additional case-specific metrics will be presented in the corresponding experimental sections. Detailed mathematical definitions of all metrics are provided in App. E.1.
@@ 4.2 Performances across different flow field solutions
To assess the model’s generalization ability across diverse flow fields, we conducted experiments using Burgers’ equation with varying initial conditions, as well as Poisson and Helmholtz equations with different analytical solutions (refer to App. C.3 and C.4 for more detailed PDE configurations). All simulations were performed on a uniform triangular mesh spanning the domain $0,1]\times[0,1]$, comprising 1,478 elements. For validation, a high-fidelity reference solution was computed on a significantly refined mesh with 23,250 elements, ensuring precise accuracy for benchmarking purposes.
The test results are summarized in Table 1, it can be observed that the our model demonstrates significant advantages in the vast majority of cases. In the seven test cases for the Poisson equation, our model achieved optimal performance (highest ER or lowest TR) in five cases, particularly excelling with complex functions. For example, for $\sum_{i,j}\exp[-\frac{(x-x_{\mu,i})^{2}}{0.25^{2}}-\frac{(y-y_{\mu,j})^{2}}{0.25^{2}}]$, ours achieved an ER of 9% , far surpassing the comparison methods. Additionally, ours achieved complete dominance in the Helmholtz equation, securing the four highest ER out of five test cases. Although slightly inferior to M2N in the Burgers equation, ours still significantly outperformed UM2N. Moreover, in all cases, our model did not produce any mesh tangling phenomena (for the mesh tangling test on non-convex meshes, refer to App. E.2). These results validate the robustness and generalization capability of our method in mesh adaptation across different PDEs, achieving state-of-the-art performance compared to existing approaches.
@@
The mesh adaptation results for the Helmholtz equation are shown in Fig. 2 (for results on the Poisson and Burgers equation, refer to App. E.3). Our method demonstrates superior visual alignment with the target flow field while simultaneously generating meshes with excellent quality compared to alternative approaches. It is noteworthy that, although M2N and UM2N can generate qualitatively adaptive meshes, they exhibit weaker compliance with the mesh equidistribution condition compared with our method (see App. E.5). Additionally, in the present work the boundary nodes are kept fixed during adaptation; experiments that allow constrained movement of boundary nodes are reported in App. E.4.
@@ 4.3 Performance across varying mesh geometries
Figure 3: The ER for different mesh resolutions on the Helmholtz equation. For clarity, we clipped the minimum ER at -1%, even though for some methods (e.g., M2N), their adapted meshes significantly increased the solution error.
@@
Regarding computational efficiency, as illustrated in Fig. 4, we present the time required for mesh movement under varying element sizes. For each model configuration, we conducted ten repeated tests on different solutions of the Helmholtz equation and reported the average mesh movement time per trial. Compared with the MA method, the neural-based mesh movement method demonstrates significant advantages. For detailed efficiency and scalability analysis, see App. E.6.
Figure 5: Results on the cylinder flow. We present the adaptive meshes at four time slices, with the results over the entire simulation period provided in the supplementary video materials.
@@
Varying mesh shapes. Another potential variable in the simulation is the shape of mesh. To evaluate the model’s generalization capability under different mesh geometric configurations, we quantitatively conducted three distinct simulation cases: subsonic flow around a NACA0012 airfoil, cylinder flow, and the wave equation on a circular domain, with the experimental setups provided in App. C.3. The adaptive results of the airfoil mesh are shown in Fig. 6, which shows that our model effectively captures the shock wave location without introducing any invalid elements. Additionally, at the position $y=0.25$, our model obtains the minimum mean absolute error (MAE  [^4] The average absolute error between the solution on the high-resolution mesh and the solution on the adaptive (coarse) mesh at each position (time step).) in pressure coefficient result. For the cylinder flow case, neither MA nor M2N could produce valid meshes. As shown in Fig. 6, compared to UM2N, our model further reduces the prediction error of the drag coefficient and slightly decreasing the cumulative error during the solving process. The results of the wave equation are shown in Fig. 8. At any given time step, our proposed method consistently generates smooth, high-quality adaptive meshes. In contrast, other methods may produce distorted mesh elements and fail to reduce errors, although they can also generate adaptive meshes.
To further validate UGM2N’s generalization to more complex mesh topologies, we demonstrated its capability on irregular and anisotropic meshes in App. E.7, where the regularity assumption (App. B.2) may not hold. Additionally, we benchmarked all methods on 1,000 random polygonal domains using Gaussian mixture flow fields as exact solutions in App. E.8. UGM2N achieves a Positive ER Ratio of 0.807 with a stable mean ER of 13.99% ± 21.05%, decisively outperforming MA (0.110) and UM2N (0.245), both of which frequently yield negative ER. These results demonstrate that UGM2N achieves substantial error reduction and robust adaptation even in challenging scenarios.
@@
Figure 8: Mesh adaptation on moving cylinder and supersonic flow cases. In supersonic flow, UGM2N’s adaptive mesh dynamically adjusts nodes to track shock waves; in the moving cylinder case, it precisely captures both the cylinder boundary and wake flow. The UM2N results for M2N are given in App. E.9. The results are also presented in the supplementary video.
Additionally, three qualitative experiments—supersonic flow over a wedge, moving cylinder in a channel, and Tohoku tsunami with complex boundaries—are provided (first two in Fig. 8, third in supplementary materials), demonstrating effective adaptive meshes for realistic simulation scenarios.
@@ 4.4 Ablation study
Loss function To validate the effectiveness of the M-Uniform loss compared to the coordinate loss of M2N and the volume loss of UM2N, we trained models with the same architecture and the same training data but different loss functions. The supervised data was generated using the MA method. Table 9 presents the average ER on PDEs with different solutions of models trained with different loss functions on three types of equations (the equation configurations are the same as in Table 1). With only a small amount of training data, supervised learning methods using the entire mesh as input struggle to produce effective models. In contrast, our unsupervised training approach requires no adaptive meshes as supervised data, and the node patch-based training method enables effective model training even with limited data. App. C.2 analysis shows increasing training data (10,440 to 41,760 patches) improves ER by up to 43% for Poisson and Helmholtz cases, indicating richer datasets further enhance model optimization.
@@
Scaling parameter $\lambda$. To assess the sensitivity of UGM2N to the scaling parameter $\lambda$ in the M-Uniform loss, we conducted an ablation study with $\lambda=10^{\mathrm{scale}}$ ($\mathrm{scale}\in\{-1,0,1,2,3\}$, 5 runs per value). The left panel of Fig. 10 shows that the converged test loss scales linearly with $\lambda$, consistent with Eq. 8, confirming that $\lambda$ modulates loss magnitude without affecting convergence dynamics. Adaptation performance, measured by error reduction, remains stable across most PDEs (right panel of Fig. 10). Minor sensitivity in Helmholtz1, Poisson4, and Poisson7 (highlighted in red) arises from case-specific optimization challenges, not model limitations, as these cases are difficult across all baselines. Overall, $\lambda$ exerts negligible influence on convergence and adaptation quality, validating the robustness of the M-Uniform loss design.
Figure 10: Ablation study on scaling parameter $\lambda$. Left: Converged test loss vs $\lambda$ (log-log scale) with theoretical linear fit (dashed). Right: Error reduction across PDEs; error bars represent ER variability across different $\lambda$ values; red bars highlight sensitive cases.
@@ 5 Conclusion
We introduce UGM2N, an unsupervised and generalizable mesh movement network. This network removes the requirement for pre-adapted meshes while demonstrating strong generalization capabilities across various PDEs, geometric configurations, and mesh resolutions. By leveraging localized node-patch representations and a novel M-Uniform loss, our approach enforces mesh equidistribution properties comparable to Monge-Ampère-based methods—but in a more efficient, unsupervised manner. Extensive experiments demonstrate consistent performance improvements over both supervised learning baselines and traditional mesh adaptation techniques, achieving significant error reductions without mesh tangling across diverse PDEs and mesh geometries. See App. F for the discussions on limitations and broader impacts.
Acknowledgment
We appreciate the reviewers for their valuable insights and helpful comments. This research was partially supported by the National Key Research and Development Program of China (2021YFB0300101, 2023YFB3001903), the National Natural Science Foundation of China (12402349), the Natural Science Foundation of Hunan Province (2024JJ6468), and the Youth Foundation of the National University of Defense Technology (ZK2023-11).
Appendix A Monge-Ampère-based mesh adaptation
Monge-Ampère (MA)-based mesh adaptation methods leverage optimal transport theory to map gradient or Hessian information from the evolving physical solution onto a mesh density function, thereby generating highly adaptive structured or unstructured meshes that dynamically refine in critical regions. The core idea is to reformulate mesh adaptation as solving the elliptic Monge-Ampère equation:
$\operatorname*{det}(D^{2}\phi(\mathbf{x}))=\frac{m(\mathbf{x})}{m_{0}},\quad\mathbf{x}\in\Omega$ (10)
where $m(\mathbf{x})$ is the monitor function, $m_{0}$ is a normalization constant, $\phi(\mathbf{x})$ is the convex scalar potential whose gradient defines the optimal coordinate transformation, and $D^{2}\phi(\mathbf{x})$ is its Hessian matrix. This yields an optimal transport mapping that automatically refines the mesh in high-gradient or high-curvature zones while coarsening it in smooth regions.
However, solving the Monge-Ampère equation demands costly nonlinear iterations—such as Newton or fixed-point methods—that are computationally expensive, sensitive to initial guesses, and often fail to converge in complex domains (e.g., non-convex geometries or those with non-smooth boundaries). These limitations severely restrict the use of MA-based methods in real-time or large-scale adaptive simulations. The present work overcomes this bottleneck by introducing an unsupervised learning framework that accelerates the MA solution process while preserving high-quality adaptive meshes.
Appendix B Method
B.1 The proposed mesh movement method